A Ferris wheel can accommodate 55 people in 25 minutes. How many people could ride the Ferris wheel in 6 hours? What was the rate per hour?


1
Expert's answer
2022-01-10T07:43:25-0500

"25min=\dfrac{25\ min}{60 \ min/h}=\dfrac{5}{12}h"

i. How many people could ride the Ferris wheel in 6 hours? 


"\dfrac{55}{\dfrac{5}{12}}=\dfrac{x}{6}"

"x=\dfrac{55(6)}{\dfrac{5}{12}}=792"

792 people could ride the Ferris wheel in 6 hours.


ii.  What was the rate per hour?


"\dfrac{792}{6}=132"

The rate is 132 people per hour.